You want to buy a new sports car from Muscle Motors for $77,000. The contract is in the form of a 72-month annuity due at an APR of 7.2 percent, compounded monthly. What will your monthly payment be? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Monthly payment           $
Answer #1

Using financial calculator

Input: Pv = -77000

N= 72

I/Y = 7.2/12 = 0.6

Find monthly payment

PMT =

$1,320.18